OpenStudy (anonymous):

If N = {all real numbers} and subset S = {multiples of 4 less than zero}, what is S′? {multiples of 4 greater than zero} {even numbers less than zero} {odd numbers less than zero} {all real numbers excluding multiples of 4 less than zero}

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

So you have \(N=\mathbb{R}\) and \(S\subset N\) with \(S=\left\{4n~:~n\in\mathbb{Z}^-\right\}\). (\(\mathbb{Z}^-\) is meant to represent the set of negative integers). The set \(S'\) is the set of all elements not in \(S\). This leaves you with every real number that's not a multiple of 4 less than 0. (the last option).